OpenStudy (anonymous):

Use the quadratic formula to sole the equation. x^2-2x= 99

OpenStudy (anonymous):

solve*

OpenStudy (unklerhaukus):

Get your quadratic into this form \[ax^2+bx+c=0\] Then apply the quadratic formula\[x_{1,2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

I need help with getting it in standard form :x

OpenStudy (unklerhaukus):

x^2-2x= 99 take away 99 from both sides

OpenStudy (anonymous):

So from -2x?

OpenStudy (unklerhaukus):

x^2-2x = 99 taking away 99 from both sides x^2-2x-99 = 99-99 =

OpenStudy (unklerhaukus):

what does the RHS become after simplification?